OpenStudy (anonymous):

What are the x-intercept(s) of the graph of the equation y - x2 = -6x - 7?

OpenStudy (zzr0ck3r):

y = -6x-7+x^2 x^2-6x-7=0 to find x int x^2+x-7x-7=0 x(x+1)-7(x+1)=0 (x-7)(x+1)=0 x = -1 and 7

OpenStudy (zzr0ck3r):

make since?

jhonyy9 (jhonyy9):

y=x^2 -6x -7 =(x-7)(x+1) than x-7=0 x=7 and x+1=0 x=-1

OpenStudy (zzr0ck3r):

no positive 7